# Is my Data in your AI Model? Membership Inference Test with Application to Face Images

### v3

@@ 6 Why Face Recognition? Exploring its Relevance and Broader Applications
7 Real-World Deployment of MINT: Challenges and Legal Implications
7.1 Obfuscation Methods
@@
7.2 Deployment and Legal Implications
@@
8 Discussion
It is important to contextualize the results achieved in the present study. While in Sec. 5.1 we provide a comparison with the state of the art in MIAs, this is for a simple interpretation as there is no specific work that aims at detecting if given data was used for training on state-of-the-art Face Recognition Models. In the state of the art of MIAs, results are available for classification tasks with small databases such as CIFAR-10 [81] where accuracy reaches up to 0.631 [82] or Yale Face [83] where accuracy is 0.558 [28]. The complexity of state-of-the-art Face Recognition models and databases is much higher, and with more relevant real-world application. Furthermore, the results achieved with the proposed MINT approach reach up to $90\%$.
On the other hand, it is important to emphasize the context of our study and the results achieved. As explained in Sec. 3, existing MIAs and our proposed MINT operate under different environmental conditions. These differences arise from the distinct application scenarios of both technologies. While MIA is designed as an attack, MINT is conceived as an auditing tool motivated on the recent legislation and aims for user transparency. As a result, MIAs require the creation of shadow models to simulate the behavior of the original model to attack, whereas in MINT we have direct access to the original model. This leads to divergent outcomes and the potential utilization of information in MINT that is not particularly useful in MIA. For instance, the use of intermediate activations seems not to yield substantial improvements in MIAs [19, 41, 40]. Nevertheless, as demonstrated in our study, appropriately leveraging this information significantly enhances performance in MINT. This implies that ineffective methods in MIA do not necessarily perform poorly in MINT, suggesting opportunities for technology-specific advancements. As an example, exploring the use of gradients could be a promising avenue for future research in MINT.
9 Conclusions
@@
This work opens numerous avenues for future research. A direct path is to enhance the presented work by incorporating additional information like the gradients of the data. The study demonstrates the effectiveness of the MINT Model even with only $1$K training samples; a subsequent step would be to explore unsupervised training and identify patterns in the activations in an unsupervised manner, eliminating the need for developer involvement. Another future direction could involve training the MINT Model alongside the Audited Model with a joint loss function to optimize performance for both tasks simultaneously. This would necessitate active collaboration from the model developer, leading to what we might call Active Membership Inference Testing, as opposed to the Passive Membership Inference Testing presented in this work. Additional lines of research include applying this technology to other applications, such as image generation models [84, 85], or to different types of data, such as text [86], and studying MINT on widely available LLMs [87].
@@ Acknowledgement
This study has been supported by projects BBforTAI (PID2021-127641OB-I00 MICINN/FEDER), HumanCAIC (TED2021-131787B-I00 MICINN) and Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2). The work of D. deAlcala is supported by a FPU Fellowship (FPU21/05785) from the Spanish MIU. A. Morales is supported by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement with Universidad Autónoma de Madrid in the line of Excellence for the University Teaching Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). The work has been conducted within the ELLIS Unit Madrid.
@@
The Image Classification domain provides a more manageable environment where numerous experiments can be conducted. Moreover, training competitive models in this domain is significantly simpler compared to Face Recognition. Results in Tab. 8. The experimental protocol follows the same structure as presented in the full paper, with additional details outlined below:
  • 1.
@@
Table 8: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) trained with $80$K for different state-of-the-art architectures.
@@
Table 9: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) trained with $80$K samples for a ResNet50, varying the resolution of the input data.
@@
When we train the models ourselves, it is also possible to analyze additional factors. In this case, we will examine the effect of input image resolution on the results. As mentioned in [42], lower input image resolution negatively impacts the performance of MIAs. Here, we aim to validate this behavior in MINT. Table 9 presents MINT performance for a ResNet50 model, varying the resolution of CIFAR-100 images prior to training the Audited Model $M$.

### v4

@@ VI Why Face Recognition? Exploring its Relevance and Broader Applications
VII Real-World Deployment of MINT: Challenges and Legal Implications
VII-A Obfuscation Methods
@@
VII-B Deployment and Legal Implications
@@
VIII Discussion
It is important to contextualize the results achieved in the present study. While in Sec. V-A we provide a comparison with the state of the art in MIAs, this is for a simple interpretation as there is no specific work that aims at detecting if given data was used for training on state-of-the-art Face Recognition Models. In the state of the art of MIAs, results are available for classification tasks with small databases such as CIFAR-10 [81] where accuracy reaches up to 0.631 [82] or Yale Face [83] where accuracy is 0.558 [28]. The complexity of state-of-the-art Face Recognition models and databases is much higher, and with more relevant real-world application. Furthermore, the results achieved with the proposed MINT approach reach up to $90\%$.
On the other hand, it is important to emphasize the context of our study and the results achieved. As explained in Sec. III, existing MIAs and our proposed MINT operate under different environmental conditions. These differences arise from the distinct application scenarios of both technologies. While MIA is designed as an attack, MINT is conceived as an auditing tool motivated on the recent legislation and aims for user transparency. As a result, MIAs require the creation of shadow models to simulate the behavior of the original model to attack, whereas in MINT we have direct access to the original model. This leads to divergent outcomes and the potential utilization of information in MINT that is not particularly useful in MIA. For instance, the use of intermediate activations seems not to yield substantial improvements in MIAs [19, 41, 40]. Nevertheless, as demonstrated in our study, appropriately leveraging this information significantly enhances performance in MINT. This implies that ineffective methods in MIA do not necessarily perform poorly in MINT, suggesting opportunities for technology-specific advancements. As an example, exploring the use of gradients could be a promising avenue for future research in MINT.
IX Conclusions
@@
This work opens numerous avenues for future research. A direct path is to enhance the presented work by incorporating additional information like the gradients of the data. The study demonstrates the effectiveness of the MINT Model even with only 11K training samples; a subsequent step would be to explore unsupervised training and identify patterns in the activations in an unsupervised manner, eliminating the need for developer involvement. Another future direction could involve training the MINT Model alongside the Audited Model with a joint loss function to optimize performance for both tasks simultaneously. This would necessitate active collaboration from the model developer, leading to what we might call Active Membership Inference Testing, as opposed to the Passive Membership Inference Testing presented in this work. Additional lines of research include applying this technology to other applications, such as image generation models [84, 85], or to different types of data, such as text [86], and studying MINT on widely available LLMs [87].
@@ Acknowledgement
This study has been supported by projects BBforTAI (PID2021-127641OB-I00 MICINN/FEDER), HumanCAIC (TED2021-131787B-I00 MICINN), M2RAI (PID2024160053OB-I00 MICIU/FEDER) and Cátedra ENIA UAM-VERIDAS en IA Responsable (NextGenerationEU PRTR TSI-100927-2023-2). The work of D. deAlcala is supported by a FPU Fellowship (FPU21/05785) from the Spanish MIU. A. Morales is supported by the Madrid Government (Comunidad de Madrid-Spain) under the Multiannual Agreement with Universidad Autónoma de Madrid in the line of Excellence for the University Teaching Staff in the context of the V PRICIT (Regional Programme of Research and Technological Innovation). The work has been conducted within the ELLIS Unit Madrid.
@@
The Image Classification domain provides a more manageable environment where numerous experiments can be conducted. Moreover, training competitive models in this domain is significantly simpler compared to Face Recognition. Results in Tab. VIII. The experimental protocol follows the same structure as presented in the full paper, with additional details outlined below:
@@
TABLE VIII: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) trained with 8080K for different state-of-the-art architectures.
@@
0.98 0.94 0.94 0.99 0.99 0.99 0.94
TABLE IX: Classification accuracy for the CNN MINT Model (Conv Layer #1 as Auditable Data) trained with 8080K samples for a ResNet50, varying the resolution of the input data.
@@
When we train the models ourselves, it is also possible to analyze additional factors. In this case, we will examine the effect of input image resolution on the results. As mentioned in [42], lower input image resolution negatively impacts the performance of MIAs. Here, we aim to validate this behavior in MINT. Table IX presents MINT performance for a ResNet50 model, varying the resolution of CIFAR-100 images prior to training the Audited Model $M$.